OpenStudy (anonymous):

Let f(x)=2n and g(x)=-n-4 Find (f o g)(n) A. -2n - 4 B. 2n - 8 C. 2n - 4 D. -2n - 8

jimthompson5910 (jim_thompson5910):

(f o g)(n) = f(g(n)) f(n) = 2n f(g(n)) = 2*(g(n)) ... replace every 'n' with g(n) f(g(n)) = 2*(-n-4) ... replace every g(n) with -n-4 on the right side I'll let you distribute and simplify that out

OpenStudy (anonymous):

So is the answer D

jimthompson5910 (jim_thompson5910):

yep